# A class of $d$-dimensional regular black holes: Shadows, Thermodynamics and Gravitational collapse

### v1

@@ Abstract
This paper examines a recently introduced class of regular black holes that can form from the collapse of a polytropic star with an arbitrary polytropic index. This class has a de Sitter core and reduces to the Bardeen and Hayward black holes when the polytropic index is chosen appropriately. We demonstrate that this class of black holes is sourced by a nonlinear electrodynamics Lagrangian in $d$ dimensions and that its regularity stems from the presence of magnetic charge. We analyze the energy conditions and study the photon spheres analytically and the shadows numerically. Then, we compare our results with observations. Additionally, we present the thermodynamic properties of this class of black holes, including their temperature, entropy, and heat capacity. We also examine their thermodynamic stability. Finally, we generalize the Oppenheimer-Snyder-Datt collapse scenario to this $d$-dimensional class of black holes and study stellar collapse into them, as well as the evolution of the star’s surface, the apparent horizon, and the event horizon.
@@ 1 Introduction
The black hole (BH) singularity is associated with the divergence of curvature invariants. It is generally believed that such a singularity would not exist in a complete theory of quantum gravity [1]. According to the weak cosmic censorship conjecture, if stellar matter satisfies the null energy condition (NEC), the BH singularity remains hidden behind the event horizon, thereby preserving the predictable structure of spacetime [2]. In general relativity, BHs represent the ultimate state of classical gravitational collapse for massive stars. However, it has been proposed that quantum effects could produce a repulsive pressure that counteracts gravitational attraction and prevents the formation of a singularity at the end of collapse [1].
@@
Here, we focus on regular BHs with de Sitter cores. Hayward and Bardeen introduced two important solutions. The Bardeen solution arises from considering general relativity coupled to nonlinear electrodynamics [8, 9], while the Hayward solution is a modified Schwarzschild BH that incorporates a fundamental length scale [10]. These models exhibit de Sitter behavior near the center and asymptotically approach Schwarzschild spacetime at larger distances. A recent paper [11] presents a general regular solution with a de Sitter core within a gravitational collapse scenario. The stellar equation of state is shown to be of the polytropic type, resulting from the smooth matching of the interior and exterior geometries of the star, without imposing any a priori assumptions about the equation of state of the collapsing matter. The end product of gravitational collapse is a general static regular BH. This solution is characterized by the polytropic index of stellar matter; the Bardeen and Hayward BHs are special cases that correspond to specific polytropic indices [11].
@@
In the context of general relativity, Tangherlini was the first to attempt to obtain a $d$-dimensional solution, which is a generalization of the Schwarzschild solution to higher dimensions [14]. Myers and Perry (MP) found exact solutions for a rotating BH in $d$ dimensions [15]. Other studies of BHs in higher dimensions can be found in [16]. The generalization to a $d$-dimensional Bardeen BH has been studied in [17].
@@
Bardeen, Carter, and Hawking attempted to identify BHs as thermodynamic systems. They related changes in mass to the horizon area through a factor proportional to the surface gravity, which they called the first law of BH thermodynamics. This law is analogous to the first law of ordinary thermodynamics, with mass, surface gravity and horizon area playing the roles of energy, temperature and entropy, respectively [26]. Hawking solidified the analogy between temperature and surface gravity by showing that the creation of quantum particles leads to BH radiation [27].
Considering the BH as a thermodynamic system characterized by temperature and entropy makes studying its heat capacity crucial. Heat capacity indicates stability (instability) for positive (negative) values, and divergence signals a possible phase transition [28, 29]. Heat capacity also determines whether a BH is stable or unstable. If a BH were to absorb mass, negative heat capacity would mean its temperature decreases because the absorption rate would exceed the emission rate. Thus, the BH would grow, and the system would become unstable. However, the presence of other factors, such as magnetic charge, can stabilize the system [30].
@@
In this paper, we first introduce a general class of regular $d$-dimensional BHs with a de Sitter core. This is a generalization of the $d$-dimensional version of a general class of regular BHs proposed in [11]. We achieve this generalization by coupling the Einstein-Hilbert action to an appropriate nonlinear electrodynamics in which the magnetic charge acts as the free parameter ensuring regularity in each dimension. This extends [36], which was done for the Bardeen BH. Next, we study the energy conditions of this model, find the circular orbits, and identify the photon spheres and shadow. We perform an exact analysis of the photon spheres and a numerical analysis of the shadow for various values of the magnetic charge and spacetime dimension up to 11. Within the model, we study BH and compact object cases to determine the limits of photon sphere and shadow existence. Using a numerical approach, we also study the thermodynamic properties of the solution, including temperature, entropy, and heat capacity. Then, we examine the stability of the model and possible phase transitions. Finally, we generalize the OSD collapse scenario by smoothly joining a $d$-dimensional flat FLRW interior to a $d$-dimensional regular BH. We then study the evolution of the surface of the star, the apparent horizon and the event horizon.
The outline of this paper is as follows: Section 2 studies a general $d$-dimensional static spherically symmetric geometry and derives the corresponding Einstein equations. Section 3 introduces a family of $d$-dimensional regular BHs with a de Sitter core using an appropriate Lagrangian. We also compute the scalar curvature, horizons, extremal conditions, and energy conditions. Section 4 analyzes photon spheres using an exact approach and shadows using a numerical approach. Section 5 delves into the thermodynamics of BHs by studying surface gravity, temperature, and entropy using an exact approach. Additionally, we assess the thermodynamic stability of BHs by identifying potential phase transitions through heat capacity analysis. Section 6 presents the scenario of a star collapsing into a $d$-dimensional regular BH and studies the evolution of the star’s surface, as well as the apparent and event horizons. Section 7 provides a summary and conclusion.
@@ 4.1 Photon spheres
Figure 4: The inner and outer horizons, and photon spheres as a function of $\tilde{q}$ for different dimensions and for $n=1$ (i.e., Hayward BH), $n=3/2$ (i.e., Bardeen BH), $n=2$ and $n=5/2$ is presented. $\tilde{q}=0$ represent ST BH.
@@ 5 BH Thermodynamics
This can be expressed in terms of the dimensionless variables obtained by normalizing $m$, i.e., the tiled variables in the previous sections, as
@@ 5.1 Entropy
As shown in Fig. 6, the temperature becomes negative below a minimum horizon radius, which is physically unacceptable. This minimum radius corresponds to the extremal horizon radius $r_{\text{ext}}$, defined as the radius at which the surface gravity (66) vanishes. It is obtained by Eqs.(31) and (32) as
@@ 6 OSD collapse in $d$ dimensions
$ds^{2}=-d\tau^{2}+\left(dr^{2}+\sqrt{1-A_{d}(r)}d\tau^{2}\right)+r^{2}d\Omega_{d-2}^{2}.$ (86)
@@
$ds^{2}_{\Sigma}$ $=-d\tau^{2}+a^{2}(\tau)R_{\text{c}}d\Omega_{d-2}^{2}\hskip 102.43008pt\text{ from inside}$ (88)
@@
In 4 dimensions, the equation of state describes a polytropic fluid $p\propto\rho^{1+1/n}$ but for higher dimensions, it describes a generalized polytropic fluid, $p\propto\rho^{1+1/n_{eff}}$, with a dimension-dependent polytropic index $n_{eff}=\frac{2(d-1)}{3(d-2)}$.
@@ 6.1 OSD collapse into $d$-dimensional regular BH
$\tau$ $=\frac{2}{d-1}\bigg(-\frac{R^{\frac{d-1}{2}}}{m^{\frac{d-3}{2}}}\,{}_{2}F_{1}\left[\sigma,\sigma;1-\sigma;-m^{\frac{1}{2}(d-2)\left(\frac{3}{n}-2\right)}q^{d-2}R^{-\frac{3(d-2)}{2n}}\right]$
$+\frac{R_{0}^{\frac{d-1}{2}}}{m^{\frac{d-3}{2}}}\,_{2}F_{1}\left[\sigma,\sigma;1-\sigma;-m^{\frac{1}{2}(d-2)\left(\frac{3}{n}-2\right)}q^{d-2}R_{0}^{-\frac{3(d-2)}{2n}}\right]\bigg),$ (99)
where $_{2}F_{1}$ is the hypergeometric function, $\sigma=-\frac{n(d-1)}{3(d-2)}$, and the integration constant $R_{0}$ is chosen such that $R=R_{0}$ at $\tau=0$. For $n=1$ and $n=3/2$ in $d=4$ we will recover the OSD collapse into Hayward and Bardeen BHs, respectively, as studied in [11]. For $q=0$, the stellar surface evolution reduces to that of the ST spacetime
@@ 6.1.1 Horizons
To determine the interior event horizon, $r_{\text{{EH}}}$, we consider the outgoing null geodesics that reach the stellar surface precisely when it intersects the exterior event horizon. As mentioned earlier, outgoing null geodesics satisfy
@@
Using Eq. (92) and the chain rule $\dot{{r}}_{\text{{EH}}}=(d{r}_{\text{{EH}}}/d{R})\dot{{R}}$, equation (107) becomes:
$\frac{d{r}_{\text{{EH}}}}{d{R}}-\frac{{r}_{\text{{EH}}}}{{R}}=-\frac{R^{\frac{d-3}{2}}}{m^{\frac{d-3}{2}}}\left(\frac{q^{d-2}+m^{\left(\frac{d-2}{2}\right)\left(\frac{3}{n}-2\right)}}{R^{\frac{3(d-2)}{2n}}}+1\right)^{\frac{n(d-1)}{3(d-2)}}.$ (108)
@@
and for Schwarzschild BH, it becomes ${r}_{\text{{EH,\,Sch}}}=3R-\frac{2R^{3/2}}{\sqrt{m}}$. An important point concerns the choice of the initial radius $R_{0}$. In addition to the condition $\tilde{R}=\tilde{R}_{0}$ at $\tilde{\tau}=0$, the initial radius must be selected such that the event horizon radius remains positive at $\tilde{\tau}=0$. For collapse to a Schwarzschild BH, the horizon radius vanishes at $\tilde{R}=9/4$ and becomes positive for $\tilde{R}_{0}>9/4$. This sets a lower bound on the admissible initial radius of the star. Otherwise, the event horizon would form before beginning the OSD collapse, which is unphysical [^8] A straightforward calculation shows that the minimum initial radius $R_{0}$ required for physical BH formation with spatially flat interior geometry is determined by $1=\int_{R_{+}}^{R_{0,\text{min}}}\frac{dR}{R\sqrt{1-f(R)}}.$ (111) where $g_{tt}=g^{rr}=f(R)$. This condition arises by demanding that the interior event horizon $r_{EH}(R)$ (here, given by Eq. (108), begins growing from zero radius at the onset of collapse. The above relation applies universally to any static, spherically symmetric exterior geometry characterized by the metric function $f(R)$. For further details and the extension to the spatially closed case see [70].. Since the Schwarzschild BH naturally has the largest $R_{0}$ limit, we chose $\tilde{R}_{0}=2.5$ in Fig. 9 to satisfy this condition.

### v2

@@ Abstract
We investigate a general class of $d$-dimensional regular black holes characterized by a de Sitter core, which arises from the gravitational collapse of a polytropic star with an arbitrary polytropic index $n$. This framework generalizes the well-known Bardeen and Hayward black holes to higher dimensions and identifies nonlinear electrodynamics with a magnetic monopole charge as the physical source ensuring spacetime regularity. We analyze the geometric structure and energy conditions, demonstrating that while the Weak and Null Energy Conditions are satisfied, the Strong Energy Condition is violated—a necessary feature for singularity avoidance. Our study of optical properties reveals the existence of stable and unstable photon spheres, with shadows persisting only up to a critical magnetic charge limit; beyond this threshold, the object becomes a horizonless compact object. Numerical results indicate that the shadow size decreases as the dimension $d$, charge $q$, or index $n$ increases, allowing for constraints based on EHT observations of M87* and SgrA*. Thermodynamically, unlike the unstable Schwarzschild-Tangherlini solution, these regular black holes exhibit regions of local stability and phase transitions, with entropy deviating from the standard area law in higher dimensions. Finally, we generalize the Oppenheimer-Snyder-Datt collapse scenario to this background. We track the evolution of horizons, the nature of the trapping horizon, and derive a critical lower bound for the initial stellar radius required for physical black hole formation. Our results show that increasing dimensions and the polytropic index delay the collapse proper time, while magnetic charge facilitates the process by reducing the minimum initial radius. These findings provide new insights into the viability of regular black holes as non-singular endpoints of gravitational collapse in higher-dimensional gravity.
@@ 1 Introduction
The black hole (BH) singularities are associated with the divergence of curvature invariants. It is generally believed that such singularities would not exist in a complete theory of quantum gravity [1]. According to the weak cosmic censorship conjecture, if stellar matter satisfies the null energy condition (NEC), the BH singularity remains hidden behind the event horizon, thereby preserving the predictable structure of spacetime [2]. In general relativity, BHs represent the ultimate state of classical gravitational collapse for massive stars. However, it has been proposed that quantum effects could produce a repulsive pressure that counteracts gravitational attraction and prevents the formation of a singularity at the end of collapse [1].
@@
In this work, we focus on regular BHs with de Sitter cores. Hayward and Bardeen introduced two important solutions. The Bardeen solution arises from considering general relativity coupled to nonlinear electrodynamics [8, 9], while the Hayward solution is a modified Schwarzschild BH that incorporates a fundamental length scale [10]. These models exhibit de Sitter behavior near the center and asymptotically approach Schwarzschild spacetime at larger distances. A recent paper [11] presents a general regular solution with a de Sitter core within a gravitational collapse scenario. The stellar equation of state is shown to be of the polytropic type, resulting from the smooth matching of the interior and exterior geometries of the star, without imposing any a priori assumptions about the equation of state of the collapsing matter. The end product of gravitational collapse is a general static regular BH. This solution is characterized by the polytropic index of stellar matter; the Bardeen and Hayward BHs are special cases that correspond to specific polytropic indices [11].
@@
In the context of general relativity, Tangherlini was the first to obtain a $d$-dimensional solution, which is a generalization of the Schwarzschild solution to higher dimensions [14]. Myers and Perry (MP) found exact solutions for a rotating BH in $d$ dimensions [15]. Other studies of BHs in higher dimensions can be found in [16]. The generalization to a $d$-dimensional Bardeen BH has been studied in [17].
@@
Bardeen, Carter, and Hawking established BHs as thermodynamic systems. They related changes in mass to the horizon area through a factor proportional to the surface gravity, which they called the first law of BH thermodynamics. This law is analogous to the first law of ordinary thermodynamics, with mass, surface gravity and horizon area playing the roles of energy, temperature and entropy, respectively [26]. Hawking solidified the analogy between temperature and surface gravity by showing that the creation of quantum particles leads to BH radiation [27].
Considering the BH as a thermodynamic system characterized by temperature and entropy makes studying its heat capacity crucial. Heat capacity indicates stability (instability) for positive (negative) values, and divergence signals a possible phase transition [28, 29]. If a BH were to absorb mass, negative heat capacity would mean its temperature decreases because the absorption rate would exceed the emission rate. Thus, the BH would grow, and the system would become unstable. However, the presence of other factors, such as magnetic charge, can stabilize the system [30].
@@
In this paper, we introduce a general class of regular $d$-dimensional BHs with a de Sitter core. This generalizes the four-dimensional regular BHs proposed in [11] to arbitrary $d$-dimensions. We achieve this generalization by coupling the Einstein-Hilbert action to an appropriate nonlinear electrodynamics theory in which the magnetic charge acts as the free parameter ensuring regularity in each dimension. This extends the work of [36], which was focused on the Bardeen BH. Next, we study the energy conditions (ECs) of this model, find the circular orbits, and identify the photon spheres and shadows. We perform an exact analysis of the photon spheres and a numerical analysis of the shadow size for various values of the magnetic charge and spacetime dimensions up to 11. Within the model, we study both BH and horizonless compact object to determine the limits of photon sphere and shadow existence. Using a numerical approach, we also study the thermodynamic properties of the solution, including temperature, entropy, and heat capacity. Then, we examine the stability of the model and possible phase transitions. Finally, we generalize the OSD collapse scenario by smoothly joining a $d$-dimensional flat FLRW interior to a $d$-dimensional regular BH. We then study the evolution of the surface of the star, the apparent horizon and the event horizon. We investigate the nature of the trapping horizon and show that increasing the number of spacetime dimensions and the polytropic index delays the collapse proper time, whereas the presence of magnetic charge facilitates the collapse scenario by reducing the minimum initial radius.
The outline of this paper is as follows: Section 2 studies a general $d$-dimensional static spherically symmetric geometry and derives the corresponding Einstein equations. Section 3 introduces a family of $d$-dimensional regular BHs with a de Sitter core using an appropriate Lagrangian. We also compute the scalar curvature, horizons, extremal conditions, and energy conditions. Section 4 analyzes photon spheres using an exact approach and shadows using a numerical approach. Section 5 delves into the thermodynamics of BHs by studying surface gravity, temperature, and entropy using analytical and numerical approaches. Additionally, we assess the thermodynamic stability of BHs by identifying potential phase transitions through heat capacity analysis. Section 6 presents the scenario of a star collapsing into a $d$-dimensional regular BH and studies the evolution of the star’s surface, as well as the apparent and event horizons. Section 7 provides a summary and conclusion.
@@ 4.1 Photon spheres
Figure 4: The inner and outer horizons, and photon spheres as a function of $\tilde{q}$ for different dimensions and for $n=1$ (i.e., Hayward BH), $n=3/2$ (i.e., Bardeen BH), $n=2$ and $n=5/2$ are presented. $\tilde{q}=0$ represents ST BH.
@@ 5 BH Thermodynamics
This can be expressed in terms of the dimensionless variables obtained by normalizing $m$, i.e., the tilde variables in the previous sections, as
@@ 5.1 Entropy
As shown in Fig. 6, the temperature becomes negative below a minimum horizon radius, which is physically unacceptable [^7] The Python code implementing the numerical methods used throughout this article is available from the authors upon request.. This minimum radius corresponds to the extremal horizon radius $r_{\text{ext}}$, defined as the radius at which the surface gravity (66) vanishes. It is obtained by Eqs.(31) and (32) as
@@ 6 OSD collapse in $d$ dimensions
$ds^{2}=-d\tau^{2}+\left(dr+\sqrt{1-A_{d}(r)}d\tau\right)^{2}+r^{2}d\Omega_{d-2}^{2}.$ (86)
@@
$ds^{2}_{\Sigma}$ $=-d\tau^{2}+a^{2}(\tau)R_{\text{c}}^{2}d\Omega_{d-2}^{2}\hskip 102.43008pt\text{ from inside}$ (88)
@@
In 4 dimensions, the equation of state describes a polytropic fluid $p\propto\rho^{1+1/n}$ but for higher dimensions, it describes a generalized polytropic fluid, $p\propto\rho^{1+1/n_{\text{eff}}}$, with a dimension-dependent polytropic index $n_{\text{eff}}=\frac{2(d-1)}{3(d-2)}$.
@@ 6.1 OSD collapse into $d$-dimensional regular BH
$\tau$ $=\frac{2}{d-1}\bigg(-\frac{R^{\frac{d-1}{2}}}{m^{\frac{d-3}{2}}}\,{}_{2}F_{1}\left[-\frac{n(d-1)}{3(d-2)},-\frac{n(d-1)}{3(d-2)};1+\frac{n(d-1)}{3(d-2)};-m^{\frac{1}{2}(d-2)\left(\frac{3}{n}-2\right)}q^{d-2}R^{-\frac{3(d-2)}{2n}}\right]$
$+\frac{R_{0}^{\frac{d-1}{2}}}{m^{\frac{d-3}{2}}}\,_{2}F_{1}\left[-\frac{n(d-1)}{3(d-2)},-\frac{n(d-1)}{3(d-2)};1+\frac{n(d-1)}{3(d-2)};-m^{\frac{1}{2}(d-2)\left(\frac{3}{n}-2\right)}q^{d-2}R_{0}^{-\frac{3(d-2)}{2n}}\right]\bigg),$ (99)
where $_{2}F_{1}$ is the hypergeometric function, and the integration constant $R_{0}$ is chosen such that $R=R_{0}$ at $\tau=0$. For $n=1$ and $n=3/2$ in $d=4$ we will recover the OSD collapse into Hayward and Bardeen BHs, respectively, as studied in [11]. For $q=0$, the stellar surface evolution reduces to that of the ST spacetime
@@ 6.1.1 Horizons
It is well-known that the union of all apparent horizons is called a trapping horizon (TH) [70] which is a hypersurface that admits a foliation by spacelike 2-surfaces on which the expansion of the outgoing null normal vanishes, $\theta_{\ell}\big|_{\text{h}}=0$, while the expansion of the ingoing null normal is nonzero, $\theta_{n}\big|_{\text{h}}\neq 0$, and $n^{a}\nabla_{a}\theta_{\ell}\big|_{\text{h}}\neq 0$. A TH is called outer if $n^{a}\nabla_{a}\theta_{\ell}\big|_{\text{h}}<0$, inner if $n^{a}\nabla_{a}\theta_{\ell}\big|_{\text{h}}>0$ , future if $\theta_{n}\big|_{\text{h}}<0$ , and past if $\theta_{n}\big|_{\text{h}}>0$. To demonstrate the THs in the collapsing scenario described above, we begin with the expansion parameter for the exterior metric [71].
$\theta_{\ell,\,\text{out}}$ $=\frac{2}{r}\left(1-\sqrt{1-A_{d}(r)}\right),$ (107)
$\theta_{n,\,\text{out}}$ $=-\frac{1}{r}\left(1+\sqrt{1-A_{d}(r)}\right).$ (108)
At the horizon, i.e., where $A_{d}(r_{\text{h}})=0$,
$\theta_{\ell,\,\text{out}}\big|_{r_{\text{h}}}=0,\hskip 28.45274pt\theta_{n,\,\text{out}}\big|_{r_{\text{h}}}=-\frac{2}{r}<0.$ (109)
This means horizons are marginally trapped tubes [^9] A marginally trapped tube [72] is a hypersurface foliated by 2-dimensional surfaces, called marginally trapped surfaces (MTS), on which the expansion of the outgoing null normal vanishes.. Furthermore,
$n^{b}\nabla_{b}\theta_{\ell,\,\text{out}}\big|_{r_{\text{h}}}=-\frac{1}{2r_{\text{h}^{2}}}(4+rA^{\prime}(r_{\text{h}})).$ (110)
In Section 5, we saw that for the larger value of $r_{\text{h}}$, i.e., the event horizon, the derivative of the metric function is always positive. Consequently, the expression above is always negative, identifying the event horizon as a future outer TH. For the interior, similar calculations lead to
$\ell^{a}_{\text{in}}$ $=(1,1+rH,0,0),$ (111)
$n^{a}_{\text{in}}$ $=\frac{1}{2}(1,-1+rH,0,0),$ (112)
and the expansion parameter can be calculated as
$\theta_{\ell,\text{in}}$ $=\frac{2}{r}\left(1+rH\right),$ (113)
$\theta_{n,\text{in}}$ $=-\frac{1}{r}\left(1-rH\right).$ (114)
For r=1/Hr=-1/H, $\theta_{\ell,\text{in}}\big|_{\text{h}}=0$, and $\theta_{n,\text{in}}\big|_{\text{h}}<0$. The Lie derivative of the expansion parameter becomes
$n^{a}\nabla_{a}\theta_{\ell}\big|_{\text{h},\text{in}}=\dot{H}+2H^{2}.$ (115)
From Friedmann’s equations $H^{2}=\frac{16\pi}{(d-1)(d-2)}\rho_{\text{star}}$ and $\dot{H}=-\frac{8\pi}{d-2}(\rho_{\text{star}}+p_{\text{star}})$, we obtain
$\dot{H}+2H^{2}=\frac{8\pi}{(d-1)(d-2)}\left[(5-d)\rho_{\text{star}}-(d-1)p_{\text{star}}\right].$ (116)
Substituting Eqs. (95) and (97), it is straightforward to verify that the resulting expression is always positive, which classifies the TH as a future inner TH. This result demonstrates that the characteristics of the TH remain unchanged in higher dimensions. As Fig. 9 shows, when the surface of the star reaches the smaller value of $r_{\text{h}}$, i.e., the Cauchy horizon, the TH disappears. Thereafter, there are no trapped surfaces inside the star. This indicates that, for a non-singular collapse scenario, the disappearance of trapped surfaces in the final stages of collapse is essential.
To determine the radius of the interior event horizon, $r_{\text{{EH}}}$, we consider the outgoing null geodesics that reach the stellar surface precisely when it intersects the exterior event horizon. As mentioned earlier, outgoing null geodesics satisfy
@@
Using Eq. (92) and the chain rule $\dot{{r}}_{\text{{EH}}}=(d{r}_{\text{{EH}}}/d{R})\dot{{R}}$, Eq. (117) becomes
$\frac{d{r}_{\text{{EH}}}}{d{R}}-\frac{{r}_{\text{{EH}}}}{{R}}$ $=-\left(1-A_{d}(R)\right)^{-1/2}$ (118)
$=-\frac{R^{\frac{d-3}{2}}}{m^{\frac{d-3}{2}}}\left(\frac{q^{d-2}+m^{\left(\frac{d-2}{2}\right)\left(\frac{3}{n}-2\right)}}{R^{\frac{3(d-2)}{2n}}}+1\right)^{\frac{n(d-1)}{3(d-2)}}.$ (119)
@@
and for Schwarzschild BH, it becomes ${r}_{\text{{EH,\,Sch}}}=3R-\frac{2R^{3/2}}{\sqrt{m}}$.
A crucial point concerns the choice of the initial radius $R_{0}$. In addition to the boundary condition ${r}_{\text{{EH}}}({r}_{\text{h}})={r}_{\text{h}}$, the initial radius of the star must be chosen such that the event horizon starts to grow from a non-negative value. This sets a lower limit for the allowed initial radius of the star. If the initial radius is smaller than this limit, the event horizon would form before the OSD collapse begins, which is unphysical. For collapse to a ST BH, the event horizon radius vanishes at $R_{0}/m=\left(\frac{d-1}{2}\right)^{\frac{2}{d-3}}$. This reduces to $R_{0,\text{min}}/m=9/4$ for the Schwarzschild BH in $4$ dimensions. Since Schwarzschild BHs naturally impose the most restrictive lower limit on $R_{0}/m$, we chose $R_{0}/m=2.5$ in Fig. 9 to satisfy this condition.
As shown in Fig. 9, let $\tau_{\text{B}}$ denote the proper time at which the event horizon begins to expand, and let the corresponding radius be $R_{\text{B}}$. By definition, $\tau_{\text{B}}=\tau(R_{\text{B}})$ and $r_{\text{EH}}(R_{\text{B}})=0$. In view of Eq. (6.1), we have
$R_{\text{B}}$ $=\frac{2R_{\text{B}}^{\frac{d-1}{2}}}{(d-3)m^{\frac{d-3}{2}}}\,{}_{2}F_{1}\left[-\frac{n(d-3)}{3(d-2)},-\frac{n(d-1)}{3(d-2)};1-\frac{n(d-3)}{3(d-2)};-\frac{m^{\left(\frac{d-2}{2}\right)\left(\frac{3}{n}-2\right)}q^{d-2}}{R_{\text{B}}^{\frac{3(d-2)}{2n}}}\right]$
$-\frac{2R_{\text{B}}{r}_{\text{h}}^{\frac{d-3}{2}}}{(d-3)m^{\frac{d-3}{2}}}\,{}_{2}F_{1}\left[-\frac{n(d-3)}{3(d-2)},-\frac{n(d-1)}{3(d-2)};1-\frac{n(d-3)}{3(d-2)};-\frac{m^{\left(\frac{d-2}{2}\right)\left(\frac{3}{n}-2\right)}q^{d-2}}{{r}_{\text{h}}^{\frac{3(d-2)}{2n}}}\right].$ (122)
The above equation clearly demonstrates that the radius $R_{\text{B}}$ is independent of the initial stellar surface radius $R_{0}$; rather, it depends only on the outer event horizon radius and the metric parameters. By employing Eq. (6.1) to determine $\tau_{\text{B}}$ and differentiating it with respect to $R_{0}$, we obtain, after straightforward simplification (see Appendix B),
$\frac{d\tau_{\text{B}}}{dR_{0}}=\frac{R_{0}^{\frac{d-3}{2}}}{m^{\frac{d-3}{2}}}\,_{2}F_{1}\left[-\frac{n(d-1)}{3(d-2)},1-\frac{n(d-1)}{3(d-2)};\frac{n(d-1)}{3(d-2)}+1;-m^{\frac{d-2}{2}\left(\frac{3}{n}-2\right)}q^{d-2}R_{0}^{-\frac{3(d-2)}{2n}}\right].$ (123)
For given values of $q$, $n$ and $d$, and for $R_{0}/m>\left(\frac{d-1}{2}\right)^{\frac{2}{d-3}}$, one can easily show that the expression above is always positive. This implies that increasing the initial stellar radius $R_{0}$ delays the formation of the event horizon, i.e., the horizon forms at a later proper time.
By solving Eq. (118) with the initial condition ${r}_{\text{{EH}}}=r_{h}$ at $R=r_{h}$, we obtain:
$r_{\text{EH}}=R\left(1-\int_{r_{\text{h}}}^{R_{0,\text{min}}}\frac{dR}{R\sqrt{1-A_{d}(R)}}\right).$ (124)
Defining the smallest possible initial radius for the star, $R_{0,\text{min}}$, as the value where $r_{\text{EH}}=0$ , we find that [^10] For further details and the extension to the spatially closed case see [73].
$1=\int_{r_{\text{h}}}^{R_{0,\text{min}}}\frac{dR}{R\sqrt{1-A_{d}(R)}}.$ (125)
Substituting Eq. (23) into Eq. (125) and then expanding it in terms of the small dimensionless magnetic charge, $\tilde{q}\ll 1$, we obtain
$1$ $\approx\int_{\tilde{r}_{\text{h}}}^{\tilde{R}_{0,\text{min}}}\tilde{R}^{\frac{d-5}{2}}d\tilde{R}+\frac{n\tilde{q}^{d-2}(d-1)}{3(d-2)}\int_{\tilde{r}_{\text{h}}}^{\tilde{R}_{0,\text{min}}}\tilde{R}^{\frac{d(n-3)-5n+6}{2n}}d\tilde{R}+\mathcal{O}(\tilde{q}^{2(d-2)})$ (126)
$=\frac{2}{d-3}\left(\tilde{R}_{0,\text{min}}^{\frac{d-3}{2}}-\tilde{r}_{\text{h}}^{\frac{d-3}{2}}\right)+\frac{2nq^{d-2}}{d(n-3)-3(n-2)}\left(\tilde{R}_{0,\text{min}}^{\frac{d(n-3)-3(n-2)}{2n}}-\tilde{r}_{\text{h}}^{\frac{d(n-3)-3(n-2)}{2n}}\right)+\mathcal{O}(\tilde{q}^{2(d-2)}).$ (127)
On the other hand, in the limit $\tilde{q}\ll 1$, the horizon radius $\tilde{r}_{\text{h}}$ can be obtained from Eq. (30)
$\tilde{r}_{\text{h}}\approx 1-\frac{2n}{3}\frac{d-1}{d-2}\tilde{q}^{d-2}+\mathcal{O}(\tilde{q}^{2(d-2)}).$ (128)
In this limit, it is expected that the minimum initial radius can be written as $\tilde{R}_{0,\text{min}}=\left(\frac{d-1}{2}\right)^{\frac{2}{d-3}}+\delta$ where $\left(\frac{d-1}{2}\right)^{\frac{2}{d-3}}$ is the ST limit (e.g., $9/4$ for $d=4$) and $\delta$ is a small correction representing the deviation of $\tilde{R}_{0,\text{min}}$ in the regular model from the ST case. To determine $\delta$, it is sufficient to substitute $\tilde{R}_{0,\text{min}}$ and (128) into Eq. (127):
$1$ $\approx\frac{2}{d-3}\left(\left(\frac{3}{2}\right)^{d-3}\left(\frac{2(d-3)}{9}\delta+1\right)-\left(1-\frac{(d-1)nq^{d-2}}{3(d-2)}\right)\right)$
$+\frac{2(d-1)}{3(d-2)}\frac{n^{2}q^{d-2}}{n(d-3)-3(d-2)}\left(\left(\frac{3}{2}\right)^{d-3-\frac{3(d-2)}{n}}-1\right)+\mathcal{O}(\tilde{q}^{2(d-2)}),$ (129)
which leads us to
$\tilde{R}_{0,\text{min}}$ $\approx\left(\frac{d-1}{2}\right)^{\frac{2}{d-3}}-\left(\frac{2}{3}\right)^{d-5}\bigg[\frac{3\left(\frac{2}{3}\right)^{4-d}-(d-1)}{d-3}$
$+\frac{2(d-1)nq^{d-2}}{3(d-2)(d-3)}+\frac{2(d-1)n^{2}q^{d-2}}{n(d-3)-3(d-2)}\left(\left(\frac{3}{2}\right)^{\frac{n(d-3)-3(d-2)}{n}}-1\right)\bigg]+\mathcal{O}(\tilde{q}^{2(d-2)}).$ (130)
The first term in the bracket vanishes at $d=4$. Moreover, its derivative with respect to $d$ is strictly positive, hence this term is positive for all $d>4$. The second term is positive for any value of $q>0$, $n>0$ and $d>4$. For the third term, consider the function $f(x)=(y^{x}-1)/x$ where $y>1$. Since $y^{x}-1$ has the same sign as $x$, the third term ($y=3/2$ and x=(n(d3)3(d2))/nx={(n(d-3)-3(d-2))}/{n}) is positive for n>3(d2)/(d3)n>3(d-2)/(d-3). If n<3(d2)/(d3)n<3(d-2)/(d-3), then both the numerator and denominator are negative. Therefore, the third term is positive in all admissible cases [^11] For the polytropic index n=3(d2)/(d3)n=3(d-2)/(d-3), one must refer back to Eq. (126) and perform the calculation directly.. Thus, the second term in Eq. (6.1.1) is always negative, which implies that small values of the magnetic charge always decrease the minimum initial radius required for the event horizon formation in the collapse scenario. In $4$ dimensions, for the Hayward case ($n=1$), $\tilde{R}_{0,\text{min}}\approx{9}/{4}-{713\tilde{q}^{2}}/{405}+\mathcal{O}(\tilde{q}^{4})$, while for the Bardeen case ($n=3/2$), $\tilde{R}_{0,\text{min}}\approx{9}/{4}-{25\tilde{q}^{2}}/{9}+\mathcal{O}(\tilde{q}^{4})$.